Given the out put Q= 2LogL + 4Logk, PL =$4, PK =$12.


Price of the goods is given Po.


A. Find the derive demand function for labour and capital.


B. Establish the supply function


C. Determine the expansion path expression for labour and capital

1
Expert's answer
2021-11-08T11:41:20-0500

Product function Q= 2logL +4log K


Deriving the input demand law for both L and K:


"\frac{\delta Q}{\delta L}=\frac{2}{Lln10}+4logK"


"\frac{\delta Q}{\delta K}=lnL+\frac{4}{Kln10}"


ย Supply function for the firm product:


Basic form of a supply function isย ๐‘„=๐‘„๐‘†(๐‘ƒ)


Q=2log4+4log12=5.52


"\therefore"ย 5.52=๐‘„"_{s}"ย (๐‘ƒ)


Q"_{s}"="\frac{5.52}{P}"


expansion path expression for labour and capital

"\frac{\delta L}{\delta K}=-\frac{PK}{PL}"


"\frac{\delta L}{\delta K}=-\frac{12}{4}"



"\frac{\\{2}{Lln10}+4logK}{\ lnL+\frac{4}{Kln10}}=-3"